A stock shows historic returns of 1.21%, 11.49%, -1.43%, .44%, and 8.93%. What is the average return over this period?

A stock shows historic returns of 7.19%, 8.32%, 13.32%, -1.52%, 3.06%, 14.96%, and 3.42%. What is the average return over this period?
Answer #1

Solution(a)

Historic returns over this period = summation(Returns)/ N = (1.21+11.49-1.43+0.44+8.93)/5 = 20.64/5 = 4.128%

Solution(b)

Historic return over this period = (7.19+8.32+13.32-1.52+3.06+14.96+3.42)/7 = 6.96%